OpenStudy (anonymous):

Equation i think 13a+7a=0 I don't know how to do this one here ... Show me?

OpenStudy (anonymous):

alright you add 13a and 7a together. so you would have 20a=0, so it would be undefined

OpenStudy (anonymous):

Oh ok lol

OpenStudy (anonymous):

Even 8z-11z=0 ?

OpenStudy (anonymous):

yeah, most likely because any fraction that has a 0 as a numerator is undefined

OpenStudy (anonymous):

K

OpenStudy (anonymous):

What about x-9-4x=-6 ?

OpenStudy (anonymous):

-4x and x= -3x, so that would be -9-3x=-6 so add 9 to both sides and you would end up with -3x=3 so divide both sides by -3 and you will end up with x=-1

OpenStudy (anonymous):

Thanks .

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What about 4c+3(c-2)=-34 ? And d+4(d+6)=-1 ?

OpenStudy (anonymous):

first distribute the 3 to (c-2) so you would have: 4c+3c-6=-34. 4c+3c would be 7c then you would have: 7c-6=-34. add 6 to both sides, then it would be 7c=-28. divide 7 by both sides and you end up with c=-4

OpenStudy (anonymous):

for the second one: distribute the 4 to d+6, then you would have: d+4d+24=-1. put like terms together, so you would have 5d+24=-1, subtract 24 from both sides and you end up with 5d=-25, divide 5 by both sides and you end up with d=-5

OpenStudy (anonymous):

So -34 n -1 will always be a negative for the final answer like d=-5?

OpenStudy (anonymous):

no if both sides are negative, it will be positive. if only one side is negative, then it will be negative